OpenStudy (jetman24):

Add the following trinomials: (5x^2 - 3x - 6) + (x^2 + 6x - 6). A. 6x^4 +3x B. 6x^2 + 3x + 12 C. 6x^2 + 3x - 12 D. 6x^2 - 12

rebeccaxhawaii (rebeccaxhawaii):

5x^2+x^2=? -3x+6x=? -6-6=?

OpenStudy (jetman24):

0

OpenStudy (jetman24):

-12

rebeccaxhawaii (rebeccaxhawaii):

 what is 1+5=?

OpenStudy (jetman24):

6

rebeccaxhawaii (rebeccaxhawaii):

 Good.